# Plan Your Travel and Travel with Your Plan: Wide-Horizon Planning and Evaluation via LLM

### v1

@@ D.2 Scaling Capability of Naive Wide-horizon Thinking
In the upper part of Figure 8, we show curves of the attention score distributions over the context (excluding the instruction). The attention scores are aggregated from the output tokens and then averaged across layers and attention heads. We observe that larger LLMs (34B and 72B) exhibit higher attention scores across the context when considering more aspects, indicating a strong ability to focus on relevant details. In contrast, the smaller LLM (7B) does not follow this pattern, achieving its highest scores when two aspects are considered. In addition, we introduce another indicator, dubbed context aware score $\mathcal{S}_{c}$, to indicate the proportion of the most attended tokens in the context out of the total, as calculated in Equation 1:
@@ E.1.1 Reward Model
Although we can leverage the results of simulation-based evaluation as the reward signals for RFT and RL training, for the sake of efficiency, we train another reward model. To be specific, we first synthesize 1K travel planning requests. We then separately use Gemini-2.0-Pro-Exp-0205 and Qwen 2.5-32B to directly generate travel plans. For about 2K generated plans, we conduct the simulation-based evaluation to get the PER scores as the reward score labels.
We conduct the pointwise generative reward modeling. We use the travel plan as the input. Instead of directly using the aggregated scalar score calculated in Equation 2, we extract the assessment text of the traveler and corresponding PER score after the whole journey as the output. We use 8 H800 GPUs to finetune the Qwen 2.5-7B model for 5 epochs with a learning rate of 2e-5, global batch size of 64.
E.1.2 Strategist Model
To train a strategist model with RFT, we construct the training dataset consisting of 50K samples with 120 distinct traveler types and 20 Chinese cities. We first conduct RFT on Deepseek-R1-Distill Qwen-7B to get the initial RFT data. We reject the sample that gets the average PER score lower than 40 after sampling $N=3$ times. We use 32 H800 GPUs to train Deepseek-Distill Qwen-7B for 3 epochs with a learning rate of 2e-5, global batch size of 128.
After training the Deepseek-R1-Distill Qwen-7B, we finally collect 22K RFT data with rejection sampling outputs. We reuse this 22K data to finetune the Qwen 2.5 7B and 32B. We use 8 H800 GPUs to train Qwen 2.5 7B for 3 epochs with a learning rate of 2e-5, global batch size of 64. We use 32 H800 GPUs to train the Qwen 2.5-32B strategist model for 3 epochs with a learning rate of 1e-5, global batch size of 128.
E.1.3 Planning Model
We first train the planning model using the 22K data above as the cold start for subsequent RL training. To train the planning model with RL, we additionally construct the training dataset consisting of 20K samples. We conduct the standard GRPO [19] using the veRL [36] framework. We use 32 H800 GPUs to train Deepseek-R1-Distill Qwen-7B, Qwen 2.5-32B, and Qwen 2.5-7B, for 200 steps with learning rates of 1e-6, 5e-7, 1e-6, respectively, and the KL loss coefficient of 0.001.
@@ E.1.4 Distillation
We reuse the input of the 22K data and use Deepseek-R1-Distill Qwen-7B (s.) + Gemini-2.5-Pro-Exp-0325 (p.) to generate the output. For each sample, we sample the output 3 times. We use the reward model to filter out the output whose reward score is lower than 80. We use the filtered dataset consisting of 15K samples to train the Qwen 2.5-3B, Llama 3.2-3B, and Deepseek-R1-Distill Qwen-7B. For the first two smaller models, we discard the thinking part. For Deepseek-R1-Distill Qwen-7B, we use the thinking part from Gemini-2.5-Pro-Exp-0325. We use 32 H800 GPUs to train Deepseek-Distill Qwen-7B for 5 epochs with a learning rate of 2e-5, global batch size of 128. We use 8 H800 GPUs to train Qwen 2.5 3B and Llama 3.2 3B for 5 epochs with a learning rate of 2e-5, global batch size of 64.
@@ E.2.1 Inference for Strategist
If we apply these aspects for subsequent analysis following the naive wide-horizon paradigm, the planning model’s performance will exhibit severe scalability limitations as the number of aspects increases. Therefore, the strategist model reorganizes these aspects and integrates them into a coherent planning blueprint. Here is the prompt used in the RFT and inference:
E.2.2 Inference for Planning Model
@@
Based on the planning blueprint, the planning model conducts a more focused and in-depth analysis following the order in the blueprint. The whole process of aspect-aware thinking is implemented through multiple turns of dialogue with shared history. We instruct the planning model to follow the guidance to analyze from one of the aspects to analyze in a turn of dialogue, as shown below:
@@ F.1.1 Action Space $\mathcal{A}$
As we have mentioned, there are four kinds of actions permitted to transit to the next state, that is, transiting, resting, dining, and sightseeing. In this section, we illustrate the details of how the travel agent takes these actions.
@@
It is noted that only resting, which means resting in place usually (If you’re heading back to the hotel to rest, choose "transit"), does not need additional processing. The following actions, including transiting, dining, and sightseeing, need to integrate with real-world information.
@@
If the traveler chooses to transit to the next location according to the plan, there are several modes of transportation, including walking, cycling, public transportation (combination of walking and bus/metro), and taxis. We use the Amap (Gaode) map API to obtain the travel time and costs of these modes of transportation. The traveler agent needs to comprehensively take into account time, cost, preferences, and stamina to make a choice.
@@ F.2 Case Study
This elderly couple is going to have a 3-day journey in Beijing. They start their journey by arriving at the Beijing West Railway Station at 10:00 AM. Their initial stamina will be subtracted by 2 ($6.5-2=4.5$) because of the exhaustion of traveling to a new city. The couple can choose one of the four actions to go to the next state. We offer four different routes utilizing various modes of transportation, all provided via the map API:
@@
Due to the limited space here, we present only a part of the simulation, but it encompasses most scenarios. This case study showcases this consistent travel simulation is based on the integration of real-world information and comprehensively reflects the quality of the generated plan.
@@ F.4 Multi-granularity Evaluation by Traveler
For every end of the visiting the POI, we let the traveler assess the POI visiting experience as follows:
@@
F.5 Human Evaluation to Verify the Effectiveness of Simulation-based Evaluation
@@ G.2.1 Feasibility (FEA): Travel Plan Similarity Score
     $max\_score\leftarrow\min(m,n)$triangleright Ideal maximum score
@@
     $completeness\_penalty\leftarrow\min(m,n)/\max(m,n)$triangleright Penalize for missing/extra activities
@@
     return $final\_similarity$
@@
     return $(time\_score+location\_score)/2$
@@
     $diff\leftarrow|$time2 - time1|||| in hours
     return $\max(0,1-diff/2)$triangleright Linear decrease within 2 hours
@@ G.2.2 Personalization (PER): Aggregated Assessment from Travelers
We evaluate the personalization of the generated plan by inspecting if the travel plan meets the unique needs and preferences of the traveler agent in the simulation. As we have mentioned, our multi-granularity evaluation mechanism via traveler’s assessment in Appendix F.4, we further introduce how we aggregate the scores from multi-granularity assessment in Equation 2 as follows:
@@
As each POI has a score after sightseeing, we average the scores of each POI to be the overall POI score $\mathcal{S}_{POI\_day\_i}$ in $i^{th}$ day. We weight the $\mathcal{S}_{POI\_day\_i}$ and the score of the whole-day itinerary $\mathcal{S}_{day\_i}$ by $\beta$ and $\gamma$ as the aggregated score of $i^{th}$ day. We set $\beta$ as 0.6 and $\gamma$ as 0.4. We average the aggregated score of each day and weight it with the score of the whole travel by $\alpha_{1}$ and $\alpha_{2}$. We set $\alpha_{1}$ as 0.6 and $\alpha_{2}$ as 0.4. We normalize the $\mathcal{S}_{whole\_travel}$ to 100.0 as the final score.

### v2

@@ D.2 Scaling Capability of Naive Wide-horizon Thinking
In the upper part of Figure 8, we show curves of the attention score distributions over the context (excluding the instruction). The attention scores are aggregated from the output tokens and then averaged across layers and attention heads. We observe that larger LLMs (34B and 72B) exhibit higher attention scores across the context when considering more aspects, indicating a strong ability to focus on relevant details. In contrast, the smaller LLM (7B) does not follow this pattern, achieving its highest scores when two aspects are considered. In addition, we introduce another indicator, dubbed context awareness score $\mathcal{S}_{c}$, to indicate the proportion of the most attended tokens in the context out of the total, as calculated in Equation 2:
@@ E.1.1 Reward Model
Training Dataset
As shown in Table 9, we collect 14 diverse cities as destinations and 37 traveler types. Based on these destinations and travelers, we first synthesize 1K travel planning requests. We then separately use Gemini-2.0-Pro-Exp-0205 and Qwen 2.5-32B to directly generate travel plans. For about 2K generated plans, we conduct the simulation-based evaluation to get the PER scores as the reward score labels.
Model Structure
To avoid reward hacking, instead of directly predicting the reward (PER-agg. score), the reward model is required to generate the corresponding scores for the five criteria in the PER score and aggregate them into the final reward.
Training Setup
We conduct the pointwise generative reward modeling. We use the travel plan as the input. Instead of directly using the aggregated scalar score calculated in Equation 3, we extract the assessment text of the traveler and the corresponding PER score after the whole journey as the output. We use 8 H800 GPUs to finetune the Qwen 2.5-7B model for 5 epochs with a learning rate of 2e-5, global batch size of 64.
E.1.2 Strategist
Training Dataset
To train a strategist with RFT, we construct the training dataset consisting of 50K samples with 120 distinct traveler types and 20 Chinese cities. We first conduct RFT on Deepseek-R1-Distill Qwen-7B to get the initial RFT data. We reject the sample that gets the average PER score lower than 40 after sampling $N=3$ times. We use 32 H800 GPUs to train Deepseek-Distill Qwen-7B for 3 epochs with a learning rate of 2e-5, global batch size of 128.
Training Setup
After training the Deepseek-R1-Distill Qwen-7B, we finally collect 22K RFT data with rejection sampling outputs. We reuse this 22K data to finetune the Qwen 2.5 7B and 32B. We use 8 H800 GPUs to train Qwen 2.5 7B for 3 epochs with a learning rate of 2e-5, global batch size of 64. We use 32 H800 GPUs to train the Qwen 2.5-32B strategist for 3 epochs with a learning rate of 1e-5, global batch size of 128.
E.1.3 Planner
Training Dataset
We first train the planner using the 22K data above as the cold start for subsequent RL training. To train the planner with RL, we additionally construct the training dataset consisting of 20K samples.
Training Setup
We conduct the standard GRPO [11] using the veRL [39] framework. We use 32 H800 GPUs to train Deepseek-R1-Distill Qwen-7B, Qwen 2.5-32B, and Qwen 2.5-7B, for 200 steps with learning rates of 1e-6, 5e-7, 1e-6, respectively, and the KL loss coefficient of 0.001. We use the KL penalty to prevent severe biases. For loss calculation, we mask the aspect-aware guidance from the strategists, preventing the planners from generating the guidance by themselves.
@@ E.1.4 Distillation
Training Dataset
We reuse the input of the 22K data and use Deepseek-R1-Distill Qwen-7B (s.) + Gemini-2.5-Pro-Exp-0325 (p.) to generate the output. For each sample, we sample the output 3 times. We use the reward model to filter out the output whose reward score is lower than 80. We use the filtered dataset consisting of 15K samples to train the Qwen 2.5-3B, Llama 3.2-3B, and Deepseek-R1-Distill Qwen-7B. For the first two smaller models, we discard the thinking part. For Deepseek-R1-Distill Qwen-7B, we use the thinking part from Gemini-2.5-Pro-Exp-0325.
Training Setup
We use 32 H800 GPUs to train Deepseek-Distill Qwen-7B for 5 epochs with a learning rate of 2e-5, global batch size of 128. We use 8 H800 GPUs to train Qwen 2.5 3B and Llama 3.2 3B for 5 epochs with a learning rate of 2e-5, global batch size of 64.
@@ E.2.1 Inference for Strategist
If we apply these aspects for subsequent analysis following the naive wide-horizon paradigm, the planner’s performance will exhibit severe scalability limitations as the number of aspects increases. Therefore, the strategist reorganizes these aspects and integrates them into a coherent planning blueprint. Here is the prompt used in the RFT and inference:
E.2.2 Inference for Planner
@@
Based on the planning blueprint, the planner conducts a more focused and in-depth analysis following the order in the blueprint. The whole process of aspect-aware thinking is implemented through multiple turns of dialogue with shared history. We instruct the planner to follow the guidance to analyze one of the aspects in a turn of dialogue, as shown below:
@@ F.1.1 Action Space $\mathcal{A}$
As we have mentioned, there are four kinds of actions permitted to transit to the next state: transiting, resting, dining, and sightseeing. In this section, we illustrate the details of how the travel agent takes these actions.
@@
It is noted that only resting, which means resting in place, usually (If you’re heading back to the hotel to rest, choose "transit"), does not need additional processing. The following actions, including transiting, dining, and sightseeing, need to be integrated with real-world information.
@@
If the traveler chooses to transit to the next location according to the plan, there are several modes of transportation, including walking, cycling, public transportation (a combination of walking and bus/metro), and taxis. We use the Amap (Gaode) map API to obtain the travel time and costs of these modes of transportation. The traveler agent needs to comprehensively take into account time, cost, preferences, and stamina to make a choice.
@@ F.2 Case Study
This elderly couple is going to have a 3-day journey in Beijing. They start their journey by arriving at the Beijing West Railway Station at 10:00 AM. Their initial stamina will be subtracted by 2 (6.52=4.56.5-2=4.5) because of the exhaustion of traveling to a new city. The couple can choose one of the four actions to go to the next state. We offer four different routes utilizing various modes of transportation, all provided via the map API:
@@
Due to the limited space here, we present only a part of the simulation, but it encompasses most scenarios. This case study showcases that causal travel simulation is based on the integration of real-world information and comprehensively reflects the quality of the generated plan.
@@ F.3 Travel-Sim Dataset Card
Table 9: Datasets with City Composition.
Dataset City List City Count Traveler Type Count
Reward Model Training
Guilin, Qingdao, Luoyang, Xishuangbanna, Shenyang,
Wuhan, Nanchang, Zhengzhou, Changchun, Xianyang,
Lanzhou, Yangzhou, Chaozhou, Guangzhou
14 37
RFT/RL Training
Beijing, Chongqing, Nanjing, Chengdu, Datong,
Jingdezhen, Dalian, Hangzhou, Beihai, Lijiang,
Kunming, Taiyuan, Xianning, Shenzhen, Hong Kong,
Changsha, Shantou, Qinhuangdao, Enshi, Tianjin
20 120
Evaluation
Lhasa, Sanya, Shanghai, Harbin, Dali, Xi’an, Xiamen
7 16
@@ F.4 Multi-granularity Evaluation by Traveler
For every end of the visit to the POI, we let the traveler assess the POI visiting experience as follows:
@@
F.5 Human Verification for Simulation-based Evaluation
@@ G.2.1 Feasibility (FEA): Travel Plan Similarity Score
  $max\_score\leftarrow\min(m,n)$ $\triangleright$ Ideal maximum score
@@
  $completeness\_penalty\leftarrow\min(m,n)/\max(m,n)$ $\triangleright$ Penalize for missing/extra activities
@@
  return final_similarityfinal\_similarity
@@
  return (time_score+location_score)/2(time\_score+location\_score)/2
@@
  $diff\leftarrow|$time2 - time1|| in hours
  return $\max(0,1-diff/2)$ $\triangleright$ Linear decrease within 2 hours
@@ G.2.2 Personalization (PER): Aggregated Assessment from Travelers
We evaluate the personalization of the generated plan by inspecting whether the travel plan meets the unique needs and preferences of the traveler agent in the simulation. As we have mentioned, our multi-granularity evaluation mechanism via traveler’s assessment in Appendix F.4, we further introduce how we aggregate the scores from multi-granularity assessment in Equation 3 as follows:
@@
As each POI has a score after sightseeing, we average the scores of each POI to be the overall POI score $\mathcal{S}_{POI\_day\_i}$ in the $i^{th}$ day. We weight the $\mathcal{S}_{POI\_day\_i}$ and the score of the whole-day itinerary $\mathcal{S}_{day\_i}$ by $\beta$ and $\gamma$ as the aggregated score of $i^{th}$ day. We set $\beta$ as 0.6 and $\gamma$ as 0.4. We average the aggregated score of each day and weight it with the score of the whole travel by $\alpha_{1}$ and $\alpha_{2}$. We set $\alpha_{1}$ as 0.6 and $\alpha_{2}$ as 0.4. We normalize the $\mathcal{S}_{whole\_travel}$ to 100.0 as the final score.